# Glimpse: Generalized Local Imaging with MLPs

### v1

@@ Abstract
Deep learning is the current de facto state of the art in tomographic imaging. A common approach is to feed the result of a simple inversion, for example the backprojection, to a convolutional neural network (CNN) which then computes the reconstruction. Despite strong results on “in-distribution” test data similar to the training data, backprojection from sparse-view data delocalizes singularities, so these approaches require a large receptive field to perform well. As a consequence, they overfit to certain global structures which leads to poor generalization on out-of-distribution (OOD) samples. Moreover, their memory complexity and training time scale unfavorably with image resolution, making them impractical for application at realistic clinical resolutions, especially in 3D: a standard U-Net requires a substantial 140GB of memory and 2600 seconds per epoch on a research-grade GPU when training on 1024 $\times$ 1024 images. In this paper, we introduce Glimpse, a local processing neural network for computed tomography which reconstructs a pixel value by feeding only the measurements associated with the neighborhood of the pixel to a simple MLP. While achieving comparable or better performance with successful CNNs like the U-Net on in-distribution test data, Glimpse significantly outperforms them on OOD samples while maintaining a memory footprint almost independent of image resolution; 5GB memory suffices to train on 1024 $\times$ 1024 images. Further, we built Glimpse to be fully differentiable, which enables feats such as recovery of accurate projection angles if they are out of calibration.
@@ I Introduction
Convolutional neural networks (CNNs) have become the standard approach for tomographic image reconstruction where the goal is to recover the image from noisy (non-)linear measurements [1]. U-Net [2] has emerged as the architecture underpinning the majority of deep learning reconstruction methods, applied with great success to a variety of imaging problems including computed tomography (CT) [3], magnetic resonance imaging (MRI) [4] and photoacoustic tomography [5]. Its success is often attributed to its multi-scale architecture [6].
@@
At the same time, certain aspects of multi-scale CNNs complicate their application to real problems. Despite satisfactory results on in-distribution test images similar to the training data, they often overfit specific image content resulting in poor generalization on out-of-distribution changes in image content and sensor shifts. Model-based networks struggle to address this drawback by integrating the forward and adjoint operators into multiple network layers or iterations [7, 8, 9, 10, 11, 12]. However, the required memory for CNNs directly scales with image resolution [13]. For instance, the widely used U-Net requires a substantial 140GB memory and 2600 seconds per epoch when training on 1024 $\times$ 1024 images using two Tesla A100 GPUs. This latter drawback is further exacerbated with model-based networks like learned primal-dual reconstruction (LPD) [8] which achieves strong performance but requires over 80GB memory even at a lower resolution $512\times 512$. This increased memory demand is due to the repeated application of the forward model and its adjoint in forward and backward passes of the neural network. This makes standard CNN-based pipelines impractical for real-world scenarios involving resolutions beyond $512\times 512$ or 3D problems.
@@
Figure 1: A point source image, its sinogram, and the sparse view FBP reconstruction; while the corresponding measurements for this pixel have sinusoidal support in the sinogram, this information is diffused along the FBP image. We stretched the contrast of the FBP image for a better visualization.
@@
To better understand the mechanics behind the poor generalization of U-Net-like CNNs which compute the reconstruction from filtered backprojections (FBP) [14], we designed an experiment as follows. Figure 1 shows an object with a point source, its sparse view sinogram measurements with sinusoidal support, and the FBP reconstruction supported over the entire field of view. This observation raises the question of the ideal receptive field size for CNNs like U-Net, as they would seemingly require a deep network with a large receptive field to cover all the information distributed across the FBP image in the input [15, 16].
@@
However, models with large receptive fields often overfit the specific image content in training data leading to poor generalization on out-of-distribution samples. Indeed, Figure 2 shows that while the U-Net produces good results when tested on in-distribution data similar to training data (here chest images), it performs poorly on out-of-distribution (OOD) brain images. This is primarily due to the large receptive field which causes overfitting to global features. This limitation makes CNNs like U-Net inappropriate for real-world applications, especially in domains such as medical imaging where robustness over distribution shifts and other uncertain and variable factors is paramount [17].
@@
Figure 2: Performance of U-Net [2] trained on chest images: evaluation on in-distribution test data (chest samples) and OOD brain samples shows that the large receptive field of the U-Net hinders its ability to generalize on OOD samples, with its PSNR even falling below that of FBP reconstruction.
@@
In this paper, we introduce a new deep learning imaging architecture termed Glimpse—a local processing neural network with a simple design adapted to the geometry of computed tomography. As shown in Figure 3, to recover the image intensity at a given target pixel, we use a multi-layer perception (MLP) that takes only the local sinogram measurements associated with this pixel and its neighbors. In contrast to CNNs like the U-Net, which globally process the input at once, Glimpse allows the MLP to focus exclusively on the local neighborhood of the sine curve. This localization results in robust performance, particularly when dealing with OOD data.
@@
Figure 3: Glimpse; a small MLP processes the measurements associated with the pixel $(x,y)$ and its neighbors extracted from the sinogram. This local processing network has promising performance on OOD data while being computationally efficient all due to its patch-based design.
@@
At the same time, this design makes Glimpse highly computationally efficient; it permits training on mini-batches of both pixels and objects. This flexibility leads to fast and efficient training, requiring a small, fixed amount of memory almost independent from the image resolution. This allows training Glimpse on large, realistic images like in resolution $1024\times 1024$.
@@
We built Glimpse to be fully differentiable, all the way down to the sensing and integration geometry. This differentiable framework presents several advantages over the standard CNN-based architectures. For instance, most methods for CT image recovery strongly rely on the sensor geometry information encoded in the forward operator, whether explicitly, as seen in methods like FBP [14], SART [18], LGS [7], and LPD [8] or implicitly as used in U-Net [2] when taking FBP as input. This fixed geometry is a problem when faced with uncertainties in calibration or blind inversion problems where the sensor geometry information is entirely unavailable. While such uncertainties might substantially degrade the quality of reconstructions of the standard methods [19, 20], our fully differentiable architecture allows the optimization of projection angles during training, which yields both an accurate estimation of the actual sensor geometry from data and high-quality reconstructions even in the presence of substantial noise in measurements and the absence of any prior information regarding sensor positions.
II Related Workd
@@
Uncalibrated CT Imaging. In CT imaging, the acquisition operator is often known but obtains insufficient measurements. This may occur when a reduced number of projections is used to minimize radiation exposure or shorten acquisition time (sparse view) or when only a limited cone of projection angles may be used (limited view). In certain situations, the acquisition operator is only partially known due to uncertain experimental conditions giving rise to uncalibrated imaging or blind inversion. Neglecting this uncertainty can result in a significant drop in the quality of the reconstructions [19]. To tackle this challenge, total least squares approaches have been developed, involving the perturbation of an assumed forward operator [23, 24, 25]. Recently, Gupta et al. [20] used autodifferentiation and gradient descent to estimate the uncalibrated forward operator in a self-consistent manner.
MLPs for Patches. A Multi-layer perceptron (MLP) is a fundamental neural network architecture used in a great variety of applications. Recently, vision transformers [26] and MLP-mixers [27] have shown promising performance in various computer vision tasks like image classification [28] and image restoration [29]. While a vanilla MLP lacks a good inductive bias for imaging, in particular translation equivariance—very much unlike CNNs—vision transformers and MLP-mixers restore it by processing patches instead of entire images [30]. However, these strategies require large datasets and networks to achieve performance comparable with CNNs. In our work, we leverage the patch-based inductive bias [31, 32] in the context of CT imaging, demonstrating that even a small MLP can achieve performance on par with or even surpassing that of popular image-to-image CNNs.
@@ III Computed Tomography
CT imaging [33] plays an important role in many applications including medical diagnosis [34], industrial testing [35], and security [36]. We consider 2D computed tomography where the image of interest $f({\mathbf{x}})$ with size $D\times D$ is reconstructed from the measurements of wave attenuation where ${\mathbf{x}}=(x,y)\in\mathbb{R}^{2}$ denotes the coordinate. The forward model is the Radon transform $Rf$ which computes integrals of $f({\mathbf{x}})$ along lines $L$,
@@
where ${\mathbf{A}}$ denotes the discretized Radon transform and we model the measurement noise by ${\mathbf{n}}$. The most commonly used analytical inversion method is the filtered backprojection (FBP),
@@
where ${\mathbf{f}}^{\text{FBP}}\in\mathbb{R}^{N\times N}$ is the FBP reconstruction, $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$, ${\mathbf{h}}$ is a certain high-pass filter, *** denotes the convolution and linear interpolation is used in (6) for evaluating $\tilde{{\mathbf{s}}}[x,\cdot]$ when $x$ is not an integer. As shown in Proposition 2 in Section S-2 in the supplementary information, while the Ram-Lak filter is the optimal choice for ${\mathbf{h}}$ in the case of noise-free complete measurements, it can amplify the noise in real-world noisy measurements, leading to poor reconstruction.
@@ III-A Glimpse: Local Imaging with MLPs
To recover the image ${\mathbf{f}}(x,y)$ at pixel $(x,y)$, we identify the elements in the sinogram ${\mathbf{s}}$ influenced by this pixel. As illustrated in Figure 1, the corresponding measurements for the pixel $(x,y)$ are supported along a sinusoidal curve in the sinogram; we denote them $\text{SIN}_{x,y}\in\mathbb{R}^{r}$, with elements being given as
@@
The proof is standard and given for completeness in Section S-3 in the supplementary information.
The sinusoidal portion of the (filtered) sinogram $\text{SIN}_{x,y}$ should have enough information to recover the pixel intensity $(x,y)$ as it contains all the measurements associated with this pixel. However, because the projections are corrupted with noise and only come from a subset of all possible projection angles, incorporating “contextual information” about the target pixel improves reconstruction quality. As shown in Figure 3, we exploit the spatial regularity of medical images (encoded in training data) by using the measurements which provide local information around $(x,y)$. This ensures that the model does not overfit large-scale features in the training data while maintaining low computational complexity. We thus additionally extract from the sinogram the regions associated with the neighboring pixels around $(x,y)$ and store this information in vector ${\mathbf{p}}_{x,y}$,
@@
where $K=C^{2}$ determines the number of neighboring pixels around $(x,y)$ for an odd number $C\geq 1$ and $d$ denotes the scale of the window which adjusts the receptive field. We concatenate the vectors inside ${\mathbf{p}}_{x,y}$ and make a long vector ${\mathbf{p}}_{x,y}\in\mathbb{R}^{r\times K}$. In order to recover the image at pixel $(x,y)$ from ${\mathbf{p}}_{x,y}$, we use a multi-layer perception $\text{MLP}_{\theta}:~{}\mathbb{R}^{r\times K}\to~{}\mathbb{R}$ parameterized by $\theta$,
$\hat{{\mathbf{f}}}_{x,y}=\text{MLP}_{\mathrm{\bm{\theta}}}\big{[}{\mathbf{p}}_{x,y}\big{]},$ (9)
@@
which approximates the pixel intensity $\hat{{\mathbf{f}}}_{x,y}$ from the local features around $(x,y)$. We call the proposed model Glimpse. In the following section, we describe how our implementation of Glimpse allows to compensate for calibration errors as well as to adapt to noisy measurements. We then propose a training strategy with resolution-agnostic memory usage in Section III-D. Further details for network architecture and training can be found in Section S-1 in supplementary information.
III-B Learned Sensor Geometry
The knowledge of projection angles $\{\alpha_{m}\}_{m=1}^{r}$ is a fundamental prerequisite in most CT image recovery algorithms, for example in FBP [14], SART [18], LGS [7], LPD [8]. Consequently, in an uncalibrated system where access to accurate sensor geometry information is lacking, the quality of reconstruction may suffer [19, 37]. To address this challenge, our top-down differentiable framework, Glimpse, allows optimizing the projection angles to retrieve $\{\alpha_{m}\}_{m=1}^{r}$ during training, see Section III-D. As we will show in experiments, Glimpse can accurately estimate the actual projection angles $\{\alpha_{m}\}_{m=1}^{r}$ purely from data resulting in high-quality reconstructions comparable with those achieved in an ideal calibrated system while maintaining strong OOD performance, unlike methods that do not attempt to correct the projection angles. As we jointly optimize $\{\alpha_{m}\}_{m=1}^{r}$ with other trainable parameters using end-to-end training, this additional angle estimation incurs a very modest computational cost. We refer to the model incorporating the learned sensor geometry as Glimpse (LSG), distinguishing it from the vanilla version, Glimpse (vanilla), which assumes a fixed sensor geometry.
III-C Adaptive Filtering for Noisy Measurements
@@
As shown in Section S-2 in the supplementary information, the Ram–Lak high-pass filter is the optimal filter ${\mathbf{h}}$ for the FBP reconstruction in the case of complete noise-free measurements. In real applications, however, we always encounter noisy projections from a subset of angles. The Ram–Lak filter is then suboptimal and typically degrades the reconstruction quality as it amplifies high-frequency noise. Alternative filters with lower amplitudes in high frequencies like Shepp–Logan, cosine, and Hamming have been used to mitigate the noisy measurements, but they are all ad hoc choices. It would be better to adapt ${\mathbf{h}}$ to the specifics of noise and sampling strategy in the target application. To design this task-specific filter, we let $\text{MLP}_{\mathrm{\bm{\theta}}}$ take as input the filtered sinogram $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$ and consider the filter ${\mathbf{h}}$ (in Fourier space) as trainable parameters to be optimized during training. This allows us to automatically learn a noise-adaptive filter from data, again with almost no additional computational cost.
III-D Resolution-agnostic Memory Usage in Training

### v2

@@ Abstract
Deep learning has become the state-of-the-art approach to medical tomographic imaging. A common approach is to feed the result of a simple inversion, for example the backprojection, to a (multiscale) convolutional neural network (CNN), which then computes the final reconstruction. Despite strong results on in-distribution test data similar to the training data, they overfit certain large-scale structures which leads to poor generalization on out-of-distribution (OOD) samples. Moreover, their memory complexity and training time scale unfavorably with image resolution, making them impractical for application at realistic clinical resolutions, especially in 3D. A standard U-Net requires a substantial 140GB of memory and 2600 seconds per epoch on a research-grade GPU when training on 1024 $\times$ 1024 images with batch size 64. In this paper, we introduce Glimpse, a local processing neural network for computed tomography which reconstructs a pixel value by processing only the measurements associated with the neighborhood of the pixel with a simple multi-layer perception (MLP). While achieving performance comparable to or better than successful CNNs like U-Net on in-distribution test data, Glimpse significantly outperforms them on OOD samples while maintaining a memory footprint almost independent of image resolution; 5GB memory suffices to train on 1024 $\times$ 1024 images, with each epoch requires 420 seconds. Because we built Glimpse to be fully differentiable it can also be used as a plug-in component of arbitrary deep learning architectures, enabling feats such as correction of miscalibrated projection orientations.
@@ I Introduction
Convolutional neural networks (CNNs) have become the standard approach for tomographic image reconstruction [1]. U-Net [2] has emerged as an architecture underpinning numerous deep learning reconstruction methods, applied with great success to a variety of imaging problems including computed tomography (CT) [3], magnetic resonance imaging (MRI) [4] and photoacoustic tomography [5]. Its success is often attributed to its particular multi-scale architecture [6].
@@
At the same time, certain aspects of multi-scale CNNs complicate their application to real problems. Despite good performance on in-distribution test images similar to the training data, they often overfit specific image content resulting in poor generalization to distribution shifts in image content and sensing as shown in this paper. Model-based networks attempt to address this drawback by integrating the forward and adjoint operators into multiple network layers or iterations [7, 8, 9, 10, 11, 12]. However, the required memory for CNNs directly scales with image resolution [13]. For instance, the widely used U-Net requires a substantial 140GB memory and 2600 seconds per epoch when training on 1024 $\times$ 1024 images using two Tesla A100 GPUs. This latter drawback is further exacerbated with model-based networks such as learned primal-dual (LPD) [8], which achieves strong performance but requires over 80GB memory and very long training time even at a lower resolution of $512\times 512$. This increased memory demand is due to the repeated application of the forward model and its adjoint in forward and backward passes of the neural network. This makes standard CNN-based pipelines impractical for real-world scenarios involving resolutions beyond $512\times 512$.
@@
Figure 1: A point source image, its sinogram, and the sparse view FBP reconstruction. While the corresponding measurements for this pixel have sinusoidal support in the sinogram, this information is diffused all over the FBP image. The contrast of the FBP image has been stretched to emphasize this effect.
@@
To better understand the mechanics behind the poor generalization of U-Net-like CNNs which compute the reconstruction from filtered backprojections (FBP) [14], we designed an experiment as follows. Figure 1 shows an object with a point source, its sparse view sinogram measurements with sinusoidal support, and the FBP reconstruction. It is evident that the FBP is supported over the entire field of view. This observation raises the question of the ideal receptive field size for CNNs like U-Net: a large receptive field may be beneficial to capture all information correlated with the value of a target pixel [15, 16].
@@
However, models with large receptive fields often overfit specific image content in training data which leads to poor generalization on out-of-distribution samples [17]. Indeed, Figure 2 shows that while U-Net produces good results when tested on in-distribution data similar to training data (here chest images), it performs poorly on out-of-distribution (here brain images). This makes CNNs like U-Net problematic in domains such as medical imaging where robustness over distribution shifts and other uncertain and variable factors is of relevance [18].
@@
Figure 2: Performance of U-Net [2] trained on chest images: evaluation on in-distribution test data (chest samples) and OOD brain samples shows that the large receptive field of U-Net hinders its ability to generalize on OOD samples, with its PSNR even falling below that of FBP reconstruction.
@@
In this paper, we introduce a new deep learning imaging architecture termed Glimpse—a simple local processing neural network adapted to the geometry of computed tomography. As shown in Figure 3, to recover the image intensity at a given target pixel, we use a MLP that takes only the local sinogram measurements associated with this pixel and its neighbors. There is no backprojection step. This localization results in robust performance, particularly when dealing with OOD data.
@@
Figure 3: Glimpse; a single MLP processes the measurements associated with the pixel $(x,y)$ and its neighbors extracted from the sinogram. This local processing network has promising performance on OOD data while being computationally efficient all due to its locality.
@@
At the same time, this design makes Glimpse highly computationally efficient; it permits training on mini-batches of both pixels and objects. This flexibility leads to fast and efficient training, requiring a small, fixed amount of memory almost independent from the image resolution. This allows training Glimpse on large, realistic images in resolution $1024\times 1024$ and beyond.
@@
We built Glimpse to be fully differentiable, all the way down to the sensing and integration geometry. This has several advantages over the standard CNN-based architectures. For instance, most methods for CT image recovery strongly rely on the sensor geometry information encoded in the forward operator, whether explicitly, as seen in methods like FBP [14], SART [19], LGS [7], and LPD [8] or implicitly as used in U-Net [2] when taking FBP as input. This fixed geometry is a problem when faced with uncertainties in calibration or blind inversion problems where the sensor geometry information is entirely unavailable. While such uncertainties might degrade the quality of reconstructions of the standard methods [20, 21], our differentiable architecture allows the optimization of projection angles which can estimate the right projection angles and improve the quality of reconstructions.
II Related Work
@@
MLPs for Imaging A multi-layer perceptron (MLP) is a fundamental neural network architecture used in a great variety of applications. Recently, vision transformers [24] and MLP-mixers [25] have shown promising performance in various computer vision tasks like image classification [26] and image restoration [27]. While, unlike CNNs, a vanilla MLP lacks a good inductive bias for imaging, in particular translation equivariance, vision transformers and MLP-mixers restore it by processing patches instead of entire images [28]. However, these strategies require large datasets and networks to achieve performance comparable with CNNs. In our work, we propose a differentiable local processing network for CT imaging, demonstrating that even a small MLP can achieve performance on par with or even surpassing that of popular image-to-image CNNs.
Uncalibrated CT Imaging. In CT imaging, the acquisition operator is often known but an insufficient number of measurements is obtained. This may occur when a reduced number of projections is used to minimize radiation exposure or shorten acquisition time (sparse view) or when only a limited cone of projection angles may be used (limited view). In certain situations, the acquisition operator is only partially or approximately known. Neglecting this uncertainty can result in a significant drop in the quality of the reconstructions [20]. To tackle this challenge, total least squares approaches have been developed, involving the perturbation of an assumed forward operator [29, 30, 31]. Recently, Gupta et al. [21] used autodifferentiation and gradient descent to estimate the uncalibrated forward operator in a self-consistent manner.
@@ III Computed Tomography
CT imaging [32] plays an important role in many applications including medical diagnosis [33], industrial testing [34], and security [35]. We consider 2D computed tomography where the image of interest $f({\mathbf{x}})$ with size $D\times D$ is reconstructed from measurements of (X-ray) attenuation. The forward model is the Radon transform $Rf$ which computes integrals of $f({\mathbf{x}})$ along lines $L$,
@@
where ${\mathbf{A}}$ is the matrix of the discretized Radon transform and we model the measurement noise by ${\mathbf{n}}$. The most commonly used analytical inversion method is the filtered backprojection (FBP),
@@
where ${\mathbf{f}}^{\text{FBP}}\in\mathbb{R}^{N\times N}$ is the FBP reconstruction, $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$, ${\mathbf{h}}$ is a certain high-pass filter, * denotes the convolution and linear interpolation is used in (6) for evaluating $\tilde{{\mathbf{s}}}[x,\cdot]$ when $x$ is not an integer. As shown in Proposition 2 in Appendix -C, while the Ram–Lak filter is the optimal choice for ${\mathbf{h}}$ in the case of noise-free complete measurements, it can amplify the noise in real-world noisy measurements, leading to poor reconstruction.
@@ III-A Glimpse: Local Imaging with MLPs
To recover the image ${\mathbf{f}}(x,y)$ at location ${\mathbf{x}}=(x,y)$, we identify the elements in the sinogram ${\mathbf{s}}$ influenced by this pixel. As illustrated in Figure 1, the corresponding measurements for the pixel $(x,y)$ are supported along a sinusoidal curve in the sinogram; we denote them $\text{SIN}_{x,y}\in\mathbb{R}^{r}$, with elements being given as
@@
$Rf(\alpha,t)$ =$=\begin{cases}1,&\text{if}\ t=r\cos(\alpha+\varphi)\\ 0,&\text{otherwise},\end{cases}$
@@
The proof is standard and outlined for completeness in Appendix -D.
The sinusoidal portion of the sinogram $\text{SIN}_{x,y}$ should have enough information to recover the pixel intensity $(x,y)$ as it contains all the measurements associated with this pixel. Note however that the pixel at $(x,y)$ influences the integral over any line passing through it and thus also the parts of the sinogram corresponding to pixels on those lines. This can be loosely thought of as a consequence of non-orthogonality of the Radon transform. The above statement is thus more precisely a statement about the filtered sinogram since information is “relocalized” by the high-pass filtering step in the FBP.
This is related to the celebrated support theorems of Sigurdur Helgason, Jan Boman, and others [36, 37, 38, 39]. These theorems state that under appropriate conditions a compactly-supported image may be recovered from a compactly-supported subset of its Radon data. These results do not involve filtering explicitly, but its influence is implicit. They apply to idealized sampling and SNR conditions.
Indeed, the high-pass filtering in the FBP is derived for noiseless data and a continuum of observed angles. In reality the projections are corrupted with noise and come from a sparse subset of projection angles. We address this by 1) incorporating “contextual information” about the target pixel and 2) letting the filter be learnable to adapt it to the specifics of discretization and noise.
As shown in Figure 3, we exploit the spatial regularity of medical images (encoded in training data) by using the measurements which provide local information around $(x,y)$. This ensures that the model does not overfit large-scale features in the training data while maintaining low computational complexity. We thus additionally extract from the sinogram the regions associated with the neighboring pixels around $(x,y)$ and store this information in vector ${\mathbf{p}}_{x,y}$,
@@
where $K=C^{2}$ determines the number of neighboring pixels around $(x,y)$ for an odd number $C\geq 1$ and $d$ denotes the scale of the window which adjusts the receptive field. In order to recover the image at pixel $(x,y)$ from ${\mathbf{p}}_{x,y}$, we use a multi-layer perception $\text{MLP}_{\theta}:~{}\mathbb{R}^{r\times K}\to~{}\mathbb{R}$ parameterized by $\theta$,
$\hat{{\mathbf{f}}}(x,y)=\text{MLP}_{\mathrm{\bm{\theta}}}\big{[}{\mathbf{p}}_{x,y}\big{]},$ (9)
@@
which estimates the pixel intensity $\hat{{\mathbf{f}}}_{x,y}$ from the local features around $(x,y)$. We call the proposed model Glimpse, standing for generalized [^1] The word “generalized” emphasizes that locality is also encoded in the transform domain, not just in real space as in some of earlier work. local imaging with MLPs. In the following section, we describe how our implementation of Glimpse allows to adapt to noisy measurements. We then propose a training strategy with resolution-agnostic memory usage in Section III-C. In Appendix -B, we show how Glimpse compensates for calibration errors. Further details for network architecture and training can be found in Appendix -A.
III-B Adaptive Filtering for Noisy Measurements
@@
The Ram–Lak high-pass filter is the optimal filter ${\mathbf{h}}$ for the FBP reconstruction in the case of complete noise-free measurements; see Appendix -C for a standard demonstration. In real applications, however, we always encounter noisy projections from a subset of angles. The Ram–Lak filter is then suboptimal and typically degrades the reconstruction quality as it amplifies high-frequency noise. Alternative filters with lower amplitudes in high frequencies like Shepp–Logan, cosine, and Hamming have been used to mitigate the noisy measurements, but they are all ad hoc choices. It is advantageous to adapt ${\mathbf{h}}$ to the specifics of noise and sampling strategy in the target application. To design this task-specific filter, we let $\text{MLP}_{\mathrm{\bm{\theta}}}$ take as input the filtered sinogram $\tilde{{\mathbf{s}}}[\cdot,m]={\mathbf{s}}[\cdot,m]*{\mathbf{h}}$ and consider the filter ${\mathbf{h}}$ (in Fourier space) as trainable parameters to be optimized during training. This allows us to automatically learn a noise-adaptive filter from data, again with almost no additional computational cost.
III-C Resolution-agnostic Memory Usage in Training